Q: The sum of the ages of a boy and his mother is 48. If the mother is three more than twice the
boy’s age, find the difference in the ages of the boy and his mother? (in the topic of simultaneous equations)
I have already figured out that B(boy) + m(mother) = 48 and that m=2b+3
thanks :)

Answer by god2012(113)

As you know :
B + M = 48 --------eq.1
M = 2B + 3 --------eq.2
Substitute 2 in 1
B + 2B + 3 = 48
B = 15
M = 33
M - B = 33-15 = 18